OpenStudy (anonymous):

hey can any1 list the main and major exponent laws plz i hav a mid term tomm plzz help

OpenStudy (unklerhaukus):

\[a^n \times a^m =a^{n+m}\]\[a^n \div a^m =a^{n-m}\]\[a^0=1\]\[(a^n)^m=a^{nm}\]\[(a \times b)^n=a^n \times b^n \]\[({a \over b})^n={a^n \over b^n}\]

OpenStudy (anonymous):

thanks so much